Xe^x+1=0 find number of roots in[-1,-1/2]

Grade:12

1 Answers

Sher Mohammad IIT Delhi
10 years ago
f(x)=Xe^x+1 , f(-1)=1-1/e and f(-1/2)= 1-1/sqrt(e), both are greater than zero , imply f is never zero between [-1,-1/2], imply zero roots

sher mohammad
b.tech, iit delhi